What is the maximum coefficient of performance (COP) for a freezer that is set to maintain the cold space at -7.2° C, which is located in a kitchen that is maintained at 23.1° C?
Tc = -7.2° C
Th = 23.1° C
1
Expert's answer
2020-12-03T10:59:12-0500
"COP_{max}=\frac{T_L}{T_H-T_L}\\COP_{max}=\frac{273.15-7.2}{23.1-(-7.2)}=8.78"